Q98 / 180 · COMEDK 2024 (Afternoon Shift) · Mathematics
There are some baskets. The chances of picking a loaded basket and choosing a red coloured one is 0.2. For every 100 tries to pick one basket, 60 times a basket is either loaded or red in colour. What is the sum of the probabilities of choosing an empty basket and choosing not a red coloured one?
- Correct answerA1.2
- B0.8
- C1.3
- D0.4
Solution
Let $L$ be the event of picking a loaded basket and $R$ be the event of picking a red-coloured basket.
Given: $P(L \cap R) = 0.2$ and $P(L \cup R) = \dfrac{60}{100} = 0.6$
$P(L) + P(R) = P(L \cup R) + P(L \cap R) = 0.6 + 0.2 = 0.8$
$P(L') + P(R') = (1 - P(L)) + (1 - P(R)) = 2 - (P(L) + P(R)) = 2 - 0.8 = 1.2$
Given: $P(L \cap R) = 0.2$ and $P(L \cup R) = \dfrac{60}{100} = 0.6$
$P(L) + P(R) = P(L \cup R) + P(L \cap R) = 0.6 + 0.2 = 0.8$
$P(L') + P(R') = (1 - P(L)) + (1 - P(R)) = 2 - (P(L) + P(R)) = 2 - 0.8 = 1.2$
Hence, the correct answer is option: 'A'.